OpenStudy (anonymous):

Can someone show me how I can find the zeros of this equation? x^4 - 8x^2 - 9 = 0

OpenStudy (solomonzelman):

let x^2=a so substitute a for x^2 (and for x^4 you would substitute a^2) can you do it now? if not say so and I'll help you more.

OpenStudy (anonymous):

Sorry, I still don't get it... Can someone help me?

OpenStudy (solomonzelman):

Yes. x^4 - 8x^2 - 9 = 0 (x^2)^2 - 8 (x^2) - 9 = 0 now we will substitute the "a" a^2 - 8a - 9 = 0 Now, which method do you prefer ? completing the square OR quadratic formula ?

OpenStudy (anonymous):

Oh, okay, I understand the substituting for "a" now. I'm not sure, I have only heard of the quadratic formula, so that I guess?

OpenStudy (solomonzelman):

So the quad. form is as follows. \[\huge\color{blue}{ \frac{-b±\sqrt{b^2-4ac} }{2a} } \] now, lets plug in your values . \[\huge\color{blue}{ \frac{-(-8)±\sqrt{(-8)^2-4(1)(-9)} }{2(1)} } \] \[\huge\color{blue}{ \frac{+8±\sqrt{64-(-36)} }{2(1)} } \] \[\huge\color{blue}{ \frac{+8±\sqrt{100} }{2(1)} } \] \[\huge\color{blue}{ \frac{+8±10 }{2} } \] \[\huge\color{blue}{ +4±5 } \] so a = 9 OR -1 So far so good?

terenzreignz (terenzreignz):

Also, just to put this there, that polynomial WAS factorable @feelgoodinc in case you feel better using that method. But Zelman's method (quadratic formula) works all the time, so you'd do well to master this method too.

OpenStudy (solomonzelman):

Hello ? This is NOT your final answer (unless you want to get it wrong) !!